OpenStudy (anonymous):

2√x+8 =3√x-2 what is x? The square root is over the x+8 and x-2

OpenStudy (anonymous):

square both sides.

OpenStudy (anonymous):

\[2(x+8)^2=3(x-2)^2 \] Like this?

OpenStudy (anonymous):

No. \[ (2\sqrt{x+8})^2 = 2^2(x+8) \]

OpenStudy (anonymous):

Then what do you do , simplify the equation on the right?

OpenStudy (anonymous):

Both sides are squared. Simplify both sides. Then try to solve for \(x\).

OpenStudy (anonymous):

can you show me how to do it please?

OpenStudy (anonymous):

\[ (2\sqrt{x+8})^2 = 4(x+8) = 4x+32 \]

OpenStudy (anonymous):

You can do the other side on your own.

OpenStudy (anonymous):

x=10?

OpenStudy (anonymous):

yes